OpenStudy (anonymous):

did i do this right? 1/2 + -7/2=-6/0=-6

OpenStudy (raden):

no, a/b + c/b = (a +c)/b so, 1/2 + -7/2 = 1/2 - 7/2 = (1 - 7)/2 = -6/2 = -3

OpenStudy (anonymous):

i am just suppost to find the sum of the equation.

OpenStudy (raden):

yeah, i see. to get the sum of 2 fractions, see if the denominators already was same then just add the numerotors

OpenStudy (raden):

* numerators

OpenStudy (anonymous):

okay i see what you are saying.

OpenStudy (anonymous):

thank you

OpenStudy (raden):

you're welcome